OpenStudy (loser66):

Find a parametrization of the curve \(x^2-y^2=1\) indicating the range of allowed values of your parametrization. Please, help

OpenStudy (loser66):

@amoodarya

OpenStudy (loser66):

@Michele_Laino

OpenStudy (michele_laino):

I think that a possible parametrization can be this: \[\Large \left\{ \begin{gathered} x = \cosh \theta \hfill \\ y = \sinh \theta \hfill \\ \end{gathered} \right.,\quad \theta \in {\mathbf{R}}\]

OpenStudy (loser66):

how about x= sec t y = tan t?

OpenStudy (michele_laino):

please keep in mind that the subsequent condition, holds: \[\Large {\left( {\cosh \theta } \right)^2} - {\left( {\sinh \theta } \right)^2} = 1\]

OpenStudy (loser66):

I know hyper trigs are work. However, if I don't know, how to derive to it?

OpenStudy (michele_laino):

I prefer hyperbolic function, since, your function is an hyperbola, and it is their natural application

OpenStudy (loser66):

There must be something to restrict tan and sec, but I don't know it. Since it is nothing wrong with tan^2 -sec^2 =1

OpenStudy (michele_laino):

I think that your parametrization also works, nevertheless it is the first time that I see it

OpenStudy (loser66):

One more question: I know the form is a hyperbola. What is the link between hyperbolic function with a hyperbola?

OpenStudy (amoodarya):

note that : \[|x| \geq 1\]

OpenStudy (loser66):

@amoodarya how do you know?

OpenStudy (loser66):

Oh, I got it @amoodarya

OpenStudy (michele_laino):

the requested link is: suppose to have your hyperbola: x^2-y^2=1 then we want to compute the area of the hyperbolic sector: